OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!! Solve for x: Log(9x+5)=3

OpenStudy (anonymous):

PLEASE

OpenStudy (anonymous):

The base is 10.

OpenStudy (anonymous):

log(x)=y so 10^y=x like that

OpenStudy (anonymous):

10^3=9x+5??

OpenStudy (anonymous):

yep, then solve x

OpenStudy (anonymous):

that's where im stuck I don't know how to

OpenStudy (anonymous):

at what?

OpenStudy (anonymous):

so do I subtract 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(9x+5=1000\) is a start

OpenStudy (anonymous):

i got it! thx